### v1

# BESSTIE: A Benchmark for Sentiment and Sarcasm Classification for Varieties of English

Abstract.
Despite large language models (LLMs) being known to exhibit bias against non-mainstream varieties, there are no known labeled datasets for sentiment analysis of English. To address this gap, we introduce BESSTIE, a benchmark for sentiment and sarcasm classification for three varieties of English: Australian (en-AU), Indian (en-IN), and British (en-UK). Using web-based content from two domains, namely, Google Place reviews and Reddit comments, we collect datasets for these language varieties using two methods: location-based and topic-based filtering. Native speakers of the language varieties manually annotate the datasets with sentiment and sarcasm labels. Subsequently, we fine-tune nine large language models (LLMs) (representing a range of encoder/decoder and mono/multilingual models) on these datasets, and evaluate their performance on the two tasks. Our results reveal that the models consistently perform better on inner-circle varieties (i.e., en-AU and en-UK), with significant performance drops for en-IN, particularly in sarcasm detection. We also report challenges in cross-variety generalisation, highlighting the need for language variety-specific datasets such as ours. BESSTIE promises to be a useful evaluative benchmark for future research in equitable LLMs, specifically in terms of language varieties. The BESSTIE datasets, code, and models are currently available on request, while the paper is under review. Please email aditya.joshi@unsw.edu.au.
bias, language variety, dialectal bias, language models, sentiment classification, sarcasm classification, benchmark, natural language processing
ccs: Computing methodologies Natural language processingccs: Computing methodologies Language resources
1. Introduction
Benchmark-based evaluation (Socher et al., 2013) of large language models (LLMs) is the prevalent norm in natural language processing (NLP). Benchmarks provide labeled datasets (Wang et al., 2019; Muennighoff et al., 2023) for specific tasks so that LLMs can be evaluated against them. However, most NLP benchmarks do not contain text in language varieties such as national varieties, dialects, sociolects or creoles (Plank, 2022; Lent et al., 2024; Joshi et al., 2024). As a result, despite the superlative performance of LLMs on Standard American English, LLMs are not evaluated on other varieties of English. Knowing that LLMs can be biased against certain language varieties (as shown by  (Deas et al., 2023) for African-American English), little empirical evidence is found for their bias towards several language varieties, particularly those of English. An exception is Multi-VALUE (Ziems et al., 2023) which creates synthetic datasets of language varieties using linguistically-informed syntactic transformations and reports a degraded performance on the GLUE benchmark. However, such synthetic texts are insufficient in measuring language variety bias because language varieties are more than syntax and encompass orthography, vocabulary, and cultural pragmatics (Nguyen, 2021). As a result, labeled datasets of natural text are crucial to measure the bias in LLMs towards these varieties.
Benchmark Sent. Sarc. Eng. Var.
Cieliebak et al. (2017) (Cieliebak et al., 2017)
Wang et al. (2018) (Wang et al., 2018)
Alharbi et al. (2020) (Alharbi et al., 2020)
Abu Farha et al. (2021) (Abu Farha et al., 2021)
Elmadany et al. (2023) (Elmadany et al., 2023)
Faisal et al. (2024) (Faisal et al., 2024)
[…]
Table 1. Comparison of BESSTIE with past benchmarks for sentiment or sarcasm classification. ‘Sent.’ indicates sentiment classification, ‘Sarc.’ denotes sarcasm classification, ‘Eng.’ denotes English, and ‘Var.’ denotes language varieties. A checkmark (✓) denotes the availability of a particular feature, while a cross (✕) indicates its absence.
To address this gap, we introduce BESSTIE, A BEnchmark for Sentiment and Sarcasm classification for varieTIes of English, namely, Australian (en-AU), Indian (en-IN), and British (en-UK). BESSTIE comprises a manually labeled dataset of textual posts along with performance results of fine-tuned LLMs on the two classification tasks. The dataset consists of text samples collected from two web-based domains, namely, Google Places reviews, and Reddit comments, using two filtering methods: location-based and topic-based filtering respectively. Following that, we obtain manual annotations for boolean sentiment and sarcasm labels from native speakers of the language varieties. The novelty of the BESSTIE dataset can be seen in Table 1. BESSTIE improves not only upon popular sentiment/sarcasm-labeled datasets but also a recent dialectal benchmark, DialectBench (Faisal et al., 2024), which contains dialectal datasets for sentiment classification but for languages other than English. Therefore, the BESSTIE dataset is novel in the introduction of new language varieties (specifically, varieties of English) and, an additional task, sarcasm classification.
We then use the dataset to model both the tasks (sentiment and sarcasm classification) as binary classification problems. We evaluate nine fine-tuned LLMs, spanning encoder/decoder and monolingual/multilingual models. Our results indicate that current models exhibit a significant decrease in performance when processing text written in en-IN, an outer-circle variety, in comparison with inner-circle varieties of English [^1] Inner-circle English varieties are those spoken by people who use English as a first language. Outer-circle varieties are those spoken by bi/multilingual speakers. (en-AU and en-UK). The performance on sarcasm classification is consistently low across the three varieties, providing pointers for future work. Following that, we examine the aggregate performance for the three varieties, the two tasks and model attributes (encoder/decoder and monolingual/multilingual). We also examine cross-variety and cross-domain performance on all models highlighting the need for BESSTIE. BESSTIE evaluation provides a solid baseline to examine biases of contemporary LLMs, and will help to develop new methods to mitigate these biases.
The contributions of this work are:
  • Creation of a manually annotated dataset of user-generated content with sentiment and sarcasm labels for three varieties of English, called BESSTIE.
  • Evaluating nine LLMs on BESSTIE and analysing their performance to identify challenges arising due to language varieties.
An overview of BESSTIE is described in Figure 1. The paper describes the dataset (Section 4; referring to ‘Collection’ and ‘Annotation’ in the figure) and the evaluation methodology (Section 5; ‘Evaluation’ in the Figure) of BESSTIE. The experiment setup and results are in Sections 6 and 7 respectively, followed by error analysis in Section 8.
2. Related work
NLP Benchmarks such as GLUE (Wang et al., 2018), SuperGLUE (Wang et al., 2019), and DynaBench (Kiela et al., 2021) have played a pivotal role in evaluating LLMs on multiple NLP tasks, including sentiment classification. However, these benchmarks are limited in their ability to capture the linguistic nuances of non-standard language varieties. Efforts to address this include datasets for African languages and dialects (Muhammad et al., 2023), Arabic dialects (Elmadany et al., 2023), and CreoleVal (Lent et al., 2023) which provides datasets for 28 creole languages. More recently, DIALECTBENCH (Faisal et al., 2024) is a collection of dialectal datasets, and associated NLP models across 281 dialectal varieties for several tasks. However, the benchmark does not contain English dialectal datasets for sentiment and sarcasm classification. BESSTIE is the first benchmark for sentiment and sarcasm classification specifically for non-standard varieties of English.
Similarly, several sarcasm classification datasets for standard variety of English exist, consisting of reddit comments (Khodak et al., 2018), amazon reviews (Filatova, 2012), and tweets (Ptáček et al., 2014; Painter et al., 2022; Abercrombie and Hovy, 2016; Oprea and Magdy, 2020). However, there is a notable absence of sarcasm classification datasets that account for non-standard varieties of English. ArSarcasm-v2 (Abu Farha et al., 2021) is a dataset annotated for sarcasm, sentiment, and the dialect of each tweet for Arabic. Abu Farha and Magdy (2021) benchmarked Arabic-specific transformer models on this dataset, noting that models trained on Arabic data, including varieties of the language, tend to perform better.
We emphasise that the papers listed above approach benchmark creation in a similar manner as BESSTIE: they create an annotated dataset for a task, and test a broad range of models to highlight the potential for future work. In that sense, these benchmark papers acknowledge that creating a dataset and measuring bias is an important task achieved by papers that introduce benchmarks, without necessarily presenting a new method itself for the task. This holds true for our paper as well.
3. Ethical Considerations
The project received ethics approval from the Human Research Ethics Committee at UNSW, Sydney (reference number: iRECS6514). We treat the reviews as separate texts and do not aggregate them by user in any way. The Reddit comments and Google reviews used in this paper are from publicly available posts, accessible through official APIs. We adhere to the terms of service and policies outlined by these platforms while gathering data. We do not attempt to de-identify any users or collect aggregate information about individual users. The dataset, including sentiment and sarcasm annotations, is shared in a manner that preserves user privacy and abides by the rules and guidelines of the source platforms.
Figure 1. Methodology of BESSTIE. (a) Collection is detailed in Section 4.1, (b) Annotation is detailed in 4.2, and (c) Evaluation is detailed in Section 5.
4. Dataset Creation
We now describe parts (a) and (b) of Figure 1 which result in a manually annotated dataset for sentiment and sarcasm classification.
4.1. Data Collection
We collect textual posts from two domains: Google Places reviews and Reddit comments, using two filtering methods: location-based and topic-based respectively. Location-based filtering implies that we select reviews that were posted for locations in the three countries. Specifically, we collect reviews posted in cities in AU, IN and UK, and their corresponding ratings ($1$ to $5$ stars, where $5$ is the highest.) using the Google Places API [^2] https://developers.google.com/maps/documentation/places/web-service/overview; Accessed on 23 July 2024 from the place types, also defined by the API. The criteria for city selection are based on population thresholds specific to each country (en-AU: $20$K; en-IN:$100$K; en-UK: $50$K). We further filter the dataset to only contain reviews with closer ratings ($2$ and $4$ stars). We also filter out any non-English content using language probabilities calculated by fastText (Grave et al., 2018) word vectors, using a threshold of $0.98$. Additionally, to mitigate the risk of including reviews written by tourists rather than residents, we exclude reviews from locations designated as ‘tourist attractions’ [^3] https://developers.google.com/maps/documentation/places/web-service/supported_types; Accessed on 13 November 2024. by the Google Places API.
To create the reddit subset, we employ topic-based filtering and choose up to four subreddits per variety (en-AU: ‘melbourne’, ‘AustralianPolitics’,‘AskAnAustralian’; en-IN: ‘India’, ‘IndiaSpeaks’, ‘BollyBlindsNGossip’; en-UK: ‘England’, ‘Britain’, ‘UnitedKingdom’, ‘GreatBritishMemes’). We select these subreddits based on the understanding that they feature popular, discussions specific to a variety, making it highly likely that contributors use the appropriate language variety. For each variety, we scrape $12,000$ comments evenly across the selected subreddits, capping at $20$ comments per post and focusing on recent posts. These are then randomly sampled and standardized to $3,000$ comments per variety before manual annotation.
4.2. Annotation
We hire three annotators [^4] Two of these are the authors of the paper who were also involved in the evaluation. each for the three language varieties. The annotators assign the processed reviews and posts two labels: sentiment and sarcasm. The choice of labels given are negative, positive, and discard. We instruct the annotators to use the discard label for uninformative examples with no apparent polarity and in a few cases, for computer-generated messages. We compensate the annotators at the casual employment rate prescribed by the host institution.
Figure 2. Confusion matrix showing the overlap in variety annotations between annotators, ant-IN, and ant-UK. Rows represent the labels assigned by ant-IN, and columns represent the labels assigned by ant-UK. The principle diagonal elements indicate agreement between annotators, while off-diagonal elements highlight disagreements.
4.3. Data Quality & Statistics
Data Quality
While previous studies utilise location-based filtering to evaluate the robustness of language technologies across varieties (Blodgett et al., 2016; Sun et al., 2023), it needs to be examined if the collected text is indeed from these national varieties. To address this, we conduct an annotation exercise with en-IN and en-UK annotators, asking them to manually identify the variety of a given text. Specifically, they label a random sample of $300$ texts ($150$ from reviews and $150$ from comments) as en-AU, en-IN, en-UK, or Cannot say. Annotator agreement is measured using Cohen’s kappa ($\kappa$).
With respect to the true label (based on the location or topic), results show higher agreement with the en-IN annotator ($\kappa$ = 0.41) compared to the en-UK annotator ($\kappa$ = 0.34), indicating fair to moderate consistency. The inter-annotator agreement itself is $0.26$. Figure 2 shows that annotators find it difficult to agree on inner-circle varieties, and have the highest agreement when identifying en-IN variety, with 46 agreements. While agreement between annotators is limited for inner-circle varieties, the higher agreement for the en-IN indicates that the subset reliably represents this variety, thereby validating our collection methodology. As a result, the two subsets (reviews and comments) collectively form a good representative sample for language varieties.
Note: We also attempt to fine-tune an encoder model fine-tuned on an unlabeled dialectal corpus, to predict the language varieties. However, this was unsuccessful due to: (a) unavailability of the dataset for en-UK, (b)the dataset being dated and inaccurate in the domain of the collected data.
Variety Comment Sentiment Sarcasm
en-AU Well set out, good stock and very friendly country staff 1 0
en-AU S**t. Christ isn’t back, is he? 0 1
en-IN Good quality foods are supplied here 1 0
en-IN Coz we all have free internet. 0 1
en-UK Traditional friendly pub. Excellent beer 1 0
en-UK What a brave potatriot 0 1
Table 2. Reviews and comments from different varieties with their corresponding annotated sentiment and sarcasm labels.
Dataset Statistics
The benchmark dataset contains annotations for two tasks: sentiment and sarcasm classification, distributed across the three varieties. Table 3 provides a breakdown of the annotated dataset (consisting of google and reddit), including the number of samples in each split (Train, Validation, and Test), along with average word length and language probability. The % Positive sentiment and % Positive sarcasm indicate the proportion of samples with positive sentiment and true sarcasm respectively. These values are the same for train, validation and test set since we perform stratified sampling. We present a few example comments along with their corresponding sentiment and sarcasm labels in Table 2. The examples illustrate the diversity of language use and the challenges involved in detecting both sentiment and sarcasm across the varieties.
Variety Subset Train Valid Test % Positive Sentiment % Positive Sarcasm $P(\bar{\texttt{eng}})$ $\bar{w}$
en-AU google 946 130 270 73% 7% 0.99 63.97
reddit 1763 241 501 32% 42% 0.98 51.72
en-IN google 1648 225 469 75% 1% 0.99 44.34
reddit 1686 230 479 25% 13% 0.87 26.92
en-UK google 1817 248 517 75% 0% 0.99 72.21
reddit 1007 138 287 12% 22% 0.98 38.04
Table 3. Dataset statistics for each variety, subset, split, and label type. $\mu$ represents the average probability of English. $P(\bar{\texttt{eng}})$ is the average language probability (calculated using fastText (Grave et al., 2018) word vectors) of a review or comment being in English. $\bar{w}$ denotes average number of words in a review/comment.
5. Evaluation Methodology
With the annotated dataset in place, we now discuss the evaluation methodology used to evaluate the LLMs on the dataset.
The training strategies for encoder and decoder models are as follows. We train encoder models using cross-entropy loss, weighted on class distribution:
$\mathcal{L}_{\texttt{{Encoder}}}=-\frac{1}{N}\sum_{j=1}^{N}\sum_{k=1}^{C}w_{k}\cdot y_{j,k}\cdot\log p_{j,k}$
Here, $N$ denotes the total number of samples in the batch, $C$ represents the number of classes (2 since the tasks are binary classification), and $w_{k}$ is the weight assigned to class k, calculated based on the class distribution ($w_{\texttt{neg}}$ for the negative and $w_{\texttt{pos}}$ for the positive class). $y_{j}$ is the expected output label for the $j$th instance where $y_{j,1}=1$ if the label is 1 and $y_{j,0}=1$ if the label is 0. $p_{j,k}$ is the probability of the $j$th instance for $k$th class.
In contrast, for decoder models, we perform zero-shot instruction fine-tuning (Ouyang et al., 2022), using maximum likelihood estimation – a standard learning objective for causal language modeling (Jain et al., 2023):
$\mathcal{L}_{\texttt{{Decoder}}}=-\frac{1}{N}\sum_{j=1}^{N}\left\{\sum_{i=|\mathbf{x}^{j}|+1}^{|\mathbf{x}^{j}|+|\mathbf{t}^{j}|}\log p(\mathbf{x}_{i}^{j}|\mathbf{x}_{<i}^{j})\right\}$
Here, $\mathbf{x}_{<i}^{j}=[\mathbf{x}_{1}^{j},\ldots,\mathbf{x}_{i-1}^{j}]$ denotes the subsequence before $\mathbf{x}_{i}^{j}$ and ||$\cdot$|| ⋅ | is the number of tokens. Specifically, $|t_{j}|$ is the number of generated tokens and is set to $1$. In other words, given the text and a task-specific prompt (given in Table 4555The forms are experimentally determined using a few test examples.), the expected behaviour of the LLM is to generate either 1 (for positive) or 0 (for negative).
Task Prompt
Sentiment Classification “Generate the sentiment of the given text. Do not give an explanation.”
Sarcasm Classification “Predict if the given text is sarcastic. Do not give an explanation.”
Table 4. Task-specific prompts used to fine-tune decoder models. Both prompts restrict the model from generating an explanation.
6. Experiment Details
We conduct our experiments on a total of nine LLMs. These include six encoder models, three trained on English corpora: BERT-Large (bert(Devlin et al., 2019), RoBERTa-Large (roberta(Liu et al., 2019), ALBERT-XXL-v2 (albert(Lan et al., 2020); and three multilingual models: Multilingual-BERT-Base (mbert), Multilingual-DistilBERT-Base (mdistil(Sanh et al., 2020), XLM-RoBERTa-Large (xlm-r(Goyal et al., 2021). We also evaluate three decoder models, one trained on English corpora: Gemma2-27B-Instruct (gemma(Gemma Team et al., 2024) and two multilingual models: Mistral-Small-Instruct-2409 (mistral(Jiang et al., 2023), Qwen2.5-72B-Instruct (qwen(Yang et al., 2024). The model-specific details including architecture, language coverage of pre-training corpus, and number of parameters are described in Table 5.
While encoder models are fine-tuned with full precision, we fine-tune quantized decoder models using QLoRA (Dettmers et al., 2023) adapters, targeting all linear layers. All models are fined-tuned for 30 epochs, with a batch size of 8 and Adam optimiser. We choose an optimal learning rate by performing a grid search over the following values: 1e-5, 2e-5, and 3e-5. All experiments are performed using two NVIDIA A100 80GB GPUs.
We report our task performances on three macro-averaged F-Score, i.e., unweighted average, disregarding the imbalance in label distribution.
Model Arch. Lang. # Params
albert Encoder English 223 M
bert Encoder English 340 M
roberta Encoder English 355 M
mbert Encoder Multilingual 177 M
mdistil Encoder Multilingual 134 M
xlm-r Encoder Multilingual 355 M
[…]
mistral Decoder Multilingual 22 B
qwen Decoder Multilingual 72 B
Table 5. Details of models including architecture (Arch.), language coverage in pretraining data (Lang.), and number of trainable parameters (# Params). Here, M is millions and B is billion.
7. Results
Figure 3. Performance comparison of various models on the sentiment classification task across different English varieties (en-AU, en-IN, and en-UK).
7.1. Overall results
We first present our results on the two tasks where models are trained and evaluated on the same variety. Figure 3 and 4 describe the model performances, reported using F-Score, on the sentiment and sarcasm classification tasks respectively. While sentiment classification is performed using both google and reddit subsets, due to the absence of sarcasm labels in google subset (as shown in Table 3), sarcasm classification is conducted only on reddit subset.
Sentiment Classification
Table 6 shows that, for sentiment classification, mistral achieves the highest average performance across all varieties, with an F-Score of 0.91 on the google subset and 0.84 on the reddit subset. In contrast, qwen reports the lowest average performance across all varieties on the google subset, with an average F-Score of 0.45, while gemma exhibits the lowest average performance across all varieties on the reddit subset, with an F-Score of 0.60.
Figure 4. Performance comparison of various models on the sarcasm classification task across different English varieties (en-AU, en-IN, and en-UK).
Sarcasm Classification
Similarly, Table 7 shows that, for sarcasm classification, both mistral and mbert achieve the highest average performance across all varieties, with an F-Score of 0.68. In contrast, gemma again reports the lowest average performance across all varieties, with an average F-Score of 0.45. The lower performance on sarcasm classification as compared to sentiment classification is expected because sarcasm often relies on not just subtle linguistic cues but also localised contemporary contextual information (Abercrombie and Hovy, 2016).
Model Enco. Mono. google-Sentiment reddit-Sentiment reddit-Sarcasm
albert 0.88 0.75 0.65
bert 0.88 0.80 0.60
roberta 0.89 0.79 0.55
mbert 0.88 0.65 0.68
mdistil 0.86 0.78 0.51
xlm-r 0.77 0.73 0.62
gemma 0.79 0.60 0.45
mistral 0.91 0.84 0.68
qwen 0.45 0.82 0.52
[…]
Table 6. Model performances, averaged across all varieties, on sentiment and sarcasm classification tasks across two subsets (google and reddit). $\mu$ represents the average performance across all models. ‘Enco.’ represents encoder models. ‘Mono.’ represents monolingual models.
7.2. Impact of Domain and Models
The average results for the models highlight the need to further probe into factors influencing model performances, namely, domain, model properties and language varieties. The results in Tables 7 and 8 describe the model performances (reporting F-Score), grouped based on model properties. Finally, task performances, averaged across all models, on varieties are presented in Table 9.
google vs. reddit
Our findings (the last row $\mu$ in Table 6), indicate that models perform better on google subset, achieving an F-Score of 0.81, compared to an F-Score of 0.75 on reddit subset. This difference can be explained by the distinct writing styles of the two sources. Reviews in google are generally more formal and informative, while posts and comments from reddit often exhibit a short (as shown in Table 3), conversational tone typical of social media. Additionally, posts and comments in reddit come from forums frequented by local speakers, which means the language and expressions used are more reflective of the variety and cultural nuances.
Domain-Task Enco. Deco.
google-Sentiment 0.86 0.72
reddit-Sentiment 0.75 0.75
reddit-Sarcasm 0.60 0.55
[…]
Table 7. Task performances, averaged across all varieties, grouped based on model architectures. $\mu$ represents the average performance across all domain-task pairs.‘Enco.’ represents encoder models while ‘Deco.’ represents decoder models.
Domain-Task Mono. Mult.
google-Sentiment 0.86 0.78
reddit-Sentiment 0.73 0.76
reddit-Sarcasm 0.56 0.60
[…]
Table 8. Task performances, averaged across all varieties, grouped based on language diversity in pre-training data. ‘Mono.’ represents monolingual models, and ‘Mult.’ represents multilingual models. $\mu$ represents the average performance across all domain-task pairs.
Model properties
Table 7 shows a consistent trend where encoder models report higher performance than decoder models across different tasks. This performance gap is expected as encoder architecture is inherently better suited for sequence classification tasks, while decoder architecture is optimised for text generation tasks. Table 8 shows that, although marginal, monolingual models report a higher average performance compared to multilingual models. This performance gap is prominent on the task involving google subset, while multilingual models perform better on tasks involving reddit subset.
Domain-Task en-AU en-IN en-UK
google-Sentiment 0.94 0.64 0.86
reddit-Sentiment 0.78 0.69 0.78
reddit-Sarcasm 0.62 0.56 0.58
[…]
Table 9. Task performances, averaged across all models, on all varieties. $\mu$ represents the average performance across all domain-task pairs.
Comparison of language varieties
Finally, Table 9 shows that models, on average, perform better on both tasks when trained on en-AU variety, closely followed by performances on en-UK variety. Both en-AU and en-UK are inner-circle varieties. Conversely, models consistently perform the worst on en-IN, an outer-circle variety.
Figure 5. Cross-variety performance analysis of mistral. The figure compares three different scenarios: pre-trained (PT), in-variety fine-tuning, and cross-variety fine-tuning for sentiment and sarcasm classification across all varieties.
Figure 6. Cross-variety performance analysis of bert. The figure compares two different scenarios: in-variety fine-tuning, and cross-variety fine-tuning for sentiment and sarcasm classification across all varieties.
Figure 7. Cross-domain performance analysis of mistral and bert comparing a decoder and an encoder model.
7.3. Cross-variety/domain evaluation
Finally, we perform cross-variety/cross-domain evaluation. This refers to the scenario where the datasets used to train and test a model are from different groups (i.e., trained on en-AU, tested on en-IN, for example). The following discussion is focused on generalisation: “can models trained on one language variety generalise well to others”.
Cross-variety evaluation
We report the cross-variety results of our best-performing model, mistral. Figure 5 compares the performance of mistral across three variants: pre-trained, in-variety fine-tuning, and cross-variety fine-tuning, for the three dataset subsets using colour-coded matrices. The rows indicate the subset that the model is trained on (i.e., en-AU, en-IN, and en-UK), along with PT, indicating that the pre-trained mistral is used. Similarly, the columns indicate the test subset (i.e., en-AU, en-IN, and en-UK). The pre-trained variant of the model (the first row in each matrix) achieves high F-Score for en-AU and en-UK, for the sentiment classification task. Fine-tuning the model improves in-variety performance (shown by the principal diagonal of the matrices), with noticeable gains in F-Score on both subsets. The model performance on cross-variety evaluation remains relatively stable across different varieties of English, indicating that domain-specific data plays a minor role in influencing cross-variety generalisation for sentiment classification.
As seen earlier, the model performs significantly worse on sarcasm classification compared to sentiment classification. While in-variety evaluation shows that fine-tuning improves over the pre-trained model with substantial increases in F-Score, it negatively impacts cross-variety generalization. This means there is a notable decrease in F-Score across other varieties when the model is fine-tuned on a specific variety. This supports the idea that sarcasm classification requires the model to grasp linguistic and cultural nuances specific to a variety, making generalisation across varieties challenging. The corresponding results for bert are in Figure 6.
Cross-domain evaluation
Similarly, to assess the cross-domain robustness of our sentiment classification models, we conduct experiments with domain data from google, and reddit. Models trained on one domain are tested on the other domain, and vice-versa, for each of the three varieties of English. Figure 7 shows that models perform better when trained on in-domain data, for both mistral and bert. We observe that pre-trained mistral performs better for all varieties in comparison to results from the cross-domain experiments.
While mistral outperforms bert at both in- and cross-domain experiments, there is significant variation in cross-domain results, highlighting the challenge in generalization, given domain variance in fine-tuned data. Moreover, the direction of cross-domain experiments also affects performance. Both models perform better for reddit $\rightarrow$ google, suggesting that models may be better at transferring sentiment from a relatively informal domain. The overall difference in cross-domain performance is significantly higher for inner-circle varieties, compared to en-IN, indicating that fine-tuning with mixed domain data may be less effective for such varieties. Across varieties, however, en-UK, an inner-circle variety, performs better for google $\rightarrow$ reddit, whereas for the bert model, this is observed with the en-IN variety.
8. Error Analysis
We randomly sample 30 examples misclassified by mistral, across all domains and varieties. We hope that these error categories will be helpful for future researchers who use the BESSTIE dataset. For sentiment classification, we identify the following error categories:
  1. (1)
    Mixed Sentiment Mis-classification: The model struggles to classify text containing both positive and negative sentiment, as in ‘The outside pool has been completely redesigned and although looks good the depth of the pool has been downsized.’.
  2. (2)
    Insufficient Context: Consider the Reddit comment ‘Darwin for getting its own dot and big font on any map of the World.’. The sentiment can only be identified with the context/information from the previous exchanges in the thread.
For sarcasm classification, the error categories are:
  1. (1)
    Contextual Sarcasm: Some instances of sarcasm rely heavily on contextual or conversational cues that are not present in the text on its own. For example, in the Reddit comment, ‘Atleast Spiderman covered his face before answering.’, the sarcastic intent can only be understood within the context of the preceding discussion in the thread.
  2. (2)
    Implicit Sarcasm: Instances where sarcasm is expressed subtly, through tone or mild exaggeration, without explicit contradictions. The example, ‘Makes more sense to just bar his company from getting further contracts whilst he is running it.’, is misidentified as not sarcastic.
  3. (3)
    False Positives from Negative Connotations: Mildly negative statements are often misclassified as sarcastic due to their phrasing or sentiment. This error is identified for examples such as ‘Sorry gal, what you experienced pretty much IS the dating culture here.’.
9. Conclusion
Biases in LLMs have been reported for several language varieties. This paper presents BESSTIE, a benchmark for sentiment and sarcasm classification across three varieties of English: Australian (en-AU), Indian (en-IN), and British (en-UK). Our evaluation spans nine LLMs, six encoders and three decoders, assessed on datasets collected from Google Places reviews (google) and Reddit comments (reddit). The models perform consistently better on en-AU and en-UK (i.e., inner-circle varieties) than on en-IN (i.e., the outer-circle variety) for both tasks. Although models report high performance on sentiment classification (F-Score of 0.81 and 0.75, for google and reddit respectively), they struggle with detecting sarcasm (F-Score of 0.59 on reddit), indicating that sarcasm classification is still largely unsolved. This is supported by our cross-variety evaluation which reveals limited generalisation capability for sarcasm classification, where cultural and contextual understanding is crucial. Notably, monolingual models marginally outperform multilingual models, suggesting that language diversity in pre-training does not extend to varieties of a language. In conclusion, BESSTIE will provide a way to measure bias of LLMs towards non-standard English varieties for the two tasks. By highlighting gaps in the model performance, especially for (a) sarcasm classification in general; and (b) outer-circle varieties like en-IN for both the tasks, this benchmark will be useful to evaluate dialectal biases of language models, and contribute to research towards equitable performance across all varieties of English.
Acknowledgments
This work was funded by Google’s Research Scholar grant awarded in 2024.